# Regge trajectories and bridges between them in integrable AdS/CFT

### v2

@@ 4.2.1 How to ride the Pomeron
$\hat{M}_{1}=-\hat{M}_{2},\;\;\;\text{ i.e. }\;\;\;\hat{\mathcal{N}}_{i}=\frac{1}{2}\left(0,2\hat{M}_{1},-2\hat{M}_{1},0\right),$ (170)
@@
It is convenient to think of what happens if we take a small deviation from the intercept, i.e. $\Delta=d/2+\delta$ with small $\delta$. Because of the degenerate asymptotics, the two columns $i=1$ and $i=4$ of $Q_{a|i}$ become linearly dependent and simultaneously explode (since the determinant is normalised) as $\delta\rightarrow 0$. However, we can perform a change of basis by considering the sum and difference of the two columns, appropriately normalising them by factors of $\delta$ such that the resulting quantities are finite as $\delta\rightarrow 0$. This procedure leads to a new column (coming from the sum) where the asymptotics is powerlike, and another one (coming from the difference) where the asymptotics contains also logarithmic terms coming from the expansion of the asymptotic powers $u^{\mathcal{N}_{a}+\delta}$. It is important to remark that these logarithmic terms only affect asymptotic expansions within specific sectors at infinity, but they do not change anything in the nature of the branch points of the functions. The only branch points remain at ±$\pm 2h+i\mathbb{Z}$ and they are all quadratic.
@@
$Q_{a|i}(u)\sim B_{a|i}u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}},\;\;i=1,2,3,$ (172)
and these functions will have a regular asymptotic expansion of the form (18) or (19) in the appropriate half of the complex plane,272727 depending on whether we are studying $Q_{a|i}^{\downarrow}$ or $Q_{a|i}^{\uparrow}$. while for the fourth column instead we require an asymptotic expansion of the form
$Q_{a|4}(u)\sim u^{0}\sum_{j=0}^{\infty}\frac{F_{a,j}}{u^{j}}+\log(u)\sum_{k=0}^{\infty}\frac{G_{a,k}}{u^{k}},\;\;\;u\rightarrow\infty,$ (173)
again valid in the appropriate half of the complex plane. In order to remove residual ambiguities in mixing the first and four columns, we demand that $F_{1,0}=0$.
@@
Finally, since the basis of Q-functions has been rotated, also the gluing matrix will change form. However, we can simply exploit the fact that one of the gluing equations is very simple, i.e., at the intercept we found that, in the conventions of this section
$\tilde{\mathbf{Q}}_{13}^{\downarrow}(u)=\alpha\;\mathbf{Q}_{13}^{\uparrow}(u)+\beta\;\mathbf{Q}_{34}^{\uparrow}(u),$ (174)
for some parameters $\alpha$, $\beta$, and this subset of the gluing equations is sufficient for the convergence of the algorithm.
@@ 4.3 Leading trajectory and Pomeron eigenvalue
Figure 5: Leading Regge trajectory (or rather its image in the $sl(2)$-like grading) at two values of the coupling constant: $h=0.1$ and $0.27$. The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black dots, their shadow by empty dots. The odd-$S$ points crossed by the trajectory correspond to non-local light-ray operators. The dotted red piecewise line denotes the BFKL-type singular limit reached at zero coupling.
@@
Figure 6: The figure illustrates some aspects of the supermultiplet structure, by which the same conformal Regge trajectory is replicated a finite number of times, each shifted by half-integer steps in the Chew-Frautschi plot generated by the supercharges. In this case, we show together the trajectory interpolating the operators at the bottom of the multiplets, together with the trajectory interpolating the operators in $sl(2)$-like grading, which is shifted by an amount $(-2,-2)$. Notice that the bottom trajectory is the leading one and, as predicted by conformal Regge theory, it indeed is symmetric under $\Delta\rightarrow 3-\Delta$. The shifted trajectories do not have individually this symmetry. Their images under $\Delta\rightarrow 3-\Delta$ are also part of the full Chew-Frautschi plot. In the following, we will be content with depicting one single representative of this multiplet structure. Usually we will plot the $sl(2)$-like grading representative, unless stated otherwise.
@@
$\frac{d^{2}\gamma(S)}{dS^{2}}\Big{|}_{\text{BPS point}}$ $=2\times 4\times\left(-\frac{7h^{2}\zeta(3)}{4}+h^{4}\left(\frac{7\pi^{2}\zeta(3)}{12}+\frac{155\zeta(5)}{8}-\frac{1}{4}\pi^{4}\log(2)\right)+\right.$ (176)
@@
where $\gamma(S)$ is the anomalous dimension on the leading trajectory.
Our numerics shows very good agreement with this prediction as well.
@@
The existence of this type of behaviour was predicted studying the analytic interpolation in spin of perturbative data in Velizhanin:2022faj , but, contrary to the case of $\mathcal{N}$=4 SYM, there is no direct diagrammatic derivation of this behaviour in this theory. However, it is known from the analysis of Velizhanin:2022faj that the leading eigenvalue is not simply related to the one of $\mathcal{N}$=4 SYM. It would be interesting to study analytically the BFKL-type regime, where we take a double scaling limit of small $\omega\equiv S_{sl(2)}+1$ and small $h$ with $h/\omega$ finite. The shape of the curve in this limit should be describable analytically with the QSC, in particular it would be interesting to generalize to ABJM the recent method of Ekhammar:2024neh to obtain several terms in the expansion. We leave this for future work.
@@ 4.4 Trajectory interpolating odd spins
It is well known that the analytic continuations interpolating even and odd values of the spin are distinct. We studied the trajectory interpolating the dimensions of the operators (175) but with $S_{sl(2)}\in 2\mathbb{N}+1$.
@@
Figure 9: Trajectory interpolating operators (175) with odd spins ($sl(2)$-like grading). The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black dots, their shadow by empty dots. The weak coupling limit (indicated by the dotted red piecewise line) is in this case a sharp intersection of two linear trajectories.
@@
Let us observe that also on the odd-spin trajectory there is a second BPS protected point, where the anomalous dimension vanishes, namely the point $\Delta_{sl(2)}=0$, $S_{sl(2)}=-1$. From the point of view of the QSC, the special nature of this point is signalled by the vanishing of some of the coefficients in front of $\mathbf{P}$ functions, indeed due to (187), we have
$\mathcal{A}_{12}\mathcal{A}_{34}(\Delta_{sl(2)}=0,S_{sl(2)}=-1)&=&0,\\ \mathcal{A}_{13}\mathcal{A}_{24}(\Delta_{sl(2)}=0,S_{sl(2)}=-1)&=&0,\\ \mathcal{A}_{14}\mathcal{A}_{23}(\Delta_{sl(2)}=0,S_{sl(2)}=-1)&=&1,\end{aligned}\quad\longleftrightarrow\quad\begin{aligned} \mathcal{A}_{12}\mathcal{A}_{34}\Big{|}_{\text{BPS point}}&=&0,\\ \mathcal{A}_{13}\mathcal{A}_{24}\Big{|}_{\text{BPS point}}&=&0,\\ \mathcal{A}_{14}\mathcal{A}_{23}\Big{|}_{\text{BPS point}}&=&1.\end{aligned}$ (177)
This is the same behaviour occurring for the BPS point living on the even-spin trajectories, i.e. the one corresponding to the stress-energy tensor, and it is the way the QSC signals a shortening condition.
The vanishing of the anomalous dimension at this point can be verified at 6 loops using the analytic continuation of the perturbative results in Velizhanin:2022faj , and is confirmed by our numerics.
@@ 4.5 Trajectories and Bridges connecting them: discovering the symmetry
$\mathcal{G}_{\text{Bridge}}\simeq(\mathcal{G}_{\text{Regge}})^{T},$ (178)
@@
$\texttt{W}_{S}:S_{\text{top}}\rightarrow-1-S_{\text{top}},\;\;\text{ i.e. }\;\;S_{\text{bottom}}\rightarrow 5-S_{\text{bottom}},\;\;\text{ i.e. }\;\;S_{sl(2)}\rightarrow 1-S_{sl(2)},$ (179)
@@
$\tau_{\text{top}}\rightarrow\tau_{\text{top}}+1+2S_{\text{top}},\;\;\;\tau_{sl(2)}\rightarrow\tau_{sl(2)}-1+2S_{sl(2)},\;\;\;\tau_{\text{bottom}}\rightarrow\tau_{\text{bottom}}-5+2S_{\text{bottom}}.$ (180)
Figure 11: This figure reproduces figure 1, but giving the $sl(2)$-like grading representatives of the trajectories in accordance with the present discussion. The coupling is $h=0.1$. The ($sl(2)$-grading copy of) the leading twist-$1$ trajectory is shown, together with two Bridging trajectories connecting it to a subleading sheet with twist $5$. The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black dots, shadows of local operators below unitarity by empty dots.
Figure 12: The panel shows, at coupling $h=0.1$ ($sl(2)$-like grading), the twist-$1$ trajectory interpolating between odd $S_{sl(2)}$ states, together with two Bridging trajectories connecting it to a subleading sheet with twist $5$. The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black dots, their shadow by empty dots.
@@
The name is explained if we follow the Bridging trajectories in the $(S,\Delta)$ plane. As we move $\Delta$, they will cross various integer values of the spin, and for each of these points we can determine whether they correspond to local operators by monitoring the gluing matrix, according to the locality criterion of section 2.3. We find that, as the spin decreases in steps of 2, they invariably reach local operators (or their spin-shadows) on subleading Regge trajectories, with twist increased in steps of $4$. At the points of intersection, we can reconstruct the full subleading trajectory by switching to the standard form of gluing matrix, $\mathcal{G}_{\text{Regge}}$. In this way, we can reach subleading sheets of the spectral surface while keeping our equations totally real.
For example, in Figure 11 is shown the Bridge shooting off from the $S_{sl(2)}=2$ operator on the leading Regge trajectory. We find that at $S_{sl(2)}=0$ (which is below the unitarity bound), it intersects a subleading Regge trajectory with twist-$5$. The Figure also shows a second Bridge between the same two trajectories, the one which intersects the local operator with $S_{sl(2)}=4$ on the leading trajectory and which reaches the same subleading trajectory at a physical local operator with $S_{sl(2)}=2$. It is obvious to guess that there are in fact infinitely many Bridges between these two Trajectories, one for each non-protected local operator on the first trajectory. We have verified this conjecture in several experiments, finding Bridges passing through every non-BPS local operator we have inspected. [^29] Notice that there is instead no Bridge passing through the BPS point. If we continue the Bridges further increasing $\Delta$ past the twist-$5$ subleading trajectory, we find that they allow us to reach further subleading sheets, with twist always increasing in steps of $4$.
We tested these numerical findings for various values of the coupling, in particular $h=1/10$ and $h=0.27$. We believe that the structure is general, and that in this way it is possible to reach infinitely many subleading Regge trajectories by using alternatively the two forms of the gluing matrix.
@@
For instance, in Figure 11 we see a Bridge passing through the $S_{sl(2)}=2$ operator on the leading twist-$1$ Regge trajectory. This can also be seen as the spin-flip of a trajectory with twist-4. The Bridge intersecting the next operator with $S_{sl(2)}=4$ is a spin-flip of a trajectory with twist-$8$, and so on: subsequent Bridges, once flipped, will be seen as trajectories of increasing twists in steps of $4$, as per (180).
[…]
The consequence of this observation is that in the Chew-Frautschi plot (which contains only Regge trajectories, not their spin-flips), we will have an unexpected symmetry relating points on different Regge trajectories. For each physical local operator sitting on a Regge trajectory of twist $\tau$, we will have an exact image, with the same anomalous dimension, which will sit below the unitarity bound on a subleading trajectory. The image is obtained by applying the Weyl reflection $\texttt{W}_{S}$, and thus where the original operator sits on a trajectory of twist $\tau$ and has co-twist $\tilde{\tau}$, its image will live on a subleading trajectory of twist $\tilde{\tau}$ and have co-twist $\tau$. On this second trajectory, the original operator (or rather its image) will appear below the unitarity bound. Thus we are learning something unexpected on the behaviour of trajectories below the unitarity bound, which is a less accessible region where the Lorentzian inversion formula does not apply. To make a very concrete example, the operator $S_{sl(2)}=2$ on the leading twist-$1$ trajectory will have an image below unitarity on a subleading trajectory with twist-$4$.

### v3

@@ 4.2.1 How to ride the Pomeron
$\hat{M}_{1}=-\hat{M}_{2},\quad\text{ i.e. }\quad\hat{\mathcal{N}}_{i}=\frac{1}{2}\left(0,2\hat{M}_{1},-2\hat{M}_{1},0\right),$ (181)
@@
It is convenient to think of what happens if we take a small deviation from the intercept, i.e. $\Delta=d/2+\epsilon$ with small $\epsilon$. Because of the degenerate asymptotics, the two columns $i=1$ and $i=4$ of $Q_{a|i}$ become linearly dependent and simultaneously explode (since the determinant is normalised) as $\epsilon\rightarrow 0$. However, we can perform a change of basis by considering the sum and difference of the two columns, normalising them appropriately by factors of $\epsilon$ such that the resulting quantities are finite as $\epsilon\rightarrow 0$. This procedure leads to a new column (coming from the sum) where the asymptotics is powerlike, and another one (coming from the difference [^30] Before taking the difference, we exploit the freedom to rescale the columns to choose $B_{a|1}=B_{a|4}$.) where the asymptotics contains also logarithmic terms coming from the expansion of the asymptotic powers $u^{\mathcal{N}_{a}+\epsilon}$. It is important to remark that these logarithmic terms only affect asymptotic expansions within specific sectors at infinity, but they do not change anything in the nature of the branch points of the functions. The only branch points remain at ±$\pm 2h+\mathrm{i}\mathbb{Z}$ and they are all quadratic.
@@
$Q_{a|i}(u)\simeq B_{a|i}u^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}},\quad i=1,2,3,$ (183)
and these functions will have a regular asymptotic expansion of the form (18) or (19) in the appropriate half of the complex plane,313131 depending on whether we are studying $Q_{a|i}^{\downarrow}$ or $Q_{a|i}^{\uparrow}$. while for the fourth column instead we require an asymptotic expansion of the form
$Q_{a|4}(u)\simeq u^{0}\sum_{j\geq 0}\frac{F_{a,j}}{u^{j}}+\log(u)\sum_{k\geq 0}\frac{G_{a,k}}{u^{k}},\quad u\rightarrow+\infty,$ (184)
again valid in the appropriate half of the complex plane. In order to remove residual ambiguities in mixing the first and fourth columns, we demand that $F_{1,0}=0$.
@@
One then proceeds imposing the gluing equations as before. The structure of the gluing matrix is formally very similar to the one of generic points on a Regge trajectory, i.e. it has a form similar to (133) (with zeros in the same places). However, we can simply exploit the fact that one of the gluing equations is very simple, i.e., at the intercept we found that, in the conventions of this section
$\tilde{\mathbf{Q}}_{13}^{\downarrow}(u)+\hat{\alpha}\,\mathbf{Q}_{13}^{\uparrow}(u)+\frac{1-\hat{\alpha}^{2}}{\hat{\beta}}\,\mathbf{Q}_{34}^{\uparrow}(u)=0,$ (185)
and this subset of the gluing equations is sufficient for the convergence of the algorithm.
@@ 4.3 Leading trajectory and Pomeron eigenvalue
Figure 5: Leading Regge trajectory (or rather its image in the $sl(2)$-like grading) at two values of the coupling constant: $h=0.1$ and $0.27$. The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black bullets, their shadow by circles. The odd-$S$ points crossed by the trajectory correspond to non-local light-ray operators. The dotted red piecewise line denotes the BFKL-type singular limit reached at zero coupling.
@@
Figure 6: The figure illustrates some aspects of the supermultiplet structure, by which the same conformal Regge trajectory is replicated a finite number of times, each shifted by half-integer steps in the Chew-Frautschi plot generated by the supercharges. In this case, we show together the trajectory interpolating the operators at the bottom of the multiplets, together with the trajectory interpolating the operators in $sl(2)$-like grading, which is shifted by an amount $(-2,-2)$. Notice that the bottom trajectory is the leading one and, as predicted by conformal Regge theory, it indeed is symmetric under $\Delta\rightarrow 3-\Delta$. The shifted trajectories do not have individually this symmetry. Their images under $\Delta\rightarrow 3-\Delta$ are also part of the full Chew-Frautschi plot. In the following, we will be content with depicting one single representative of this multiplet structure. Usually we will plot the $sl(2)$-like grading representative, unless stated otherwise. The red point marks the position of the stress energy tensor (on the bottom trajectory), or correspondingly a BPS scalar operator in the same supermultiplet (on the $sl(2)$ trajectory), i.e. (186) with $S=0$.
@@
$\frac{d^{2}\gamma(S)}{dS^{2}}\Big{|}_{\text{BPS point}}$ $=2\times 4\times\left(-\frac{7h^{2}\zeta(3)}{4}+h^{4}\left(\frac{7\pi^{2}\zeta(3)}{12}+\frac{155\zeta(5)}{8}-\frac{1}{4}\pi^{4}\log(2)\right)\right.$ (187)
@@
where $\gamma(S)$ is the anomalous dimension on the leading trajectory. Our numerics shows very good agreement with this prediction as well.
@@
The existence of this type of behaviour was predicted studying the analytic interpolation in spin of perturbative data in Velizhanin:2022faj , but, contrary to the case of $\mathcal{N}$$=$$4$ SYM, there is no direct diagrammatic derivation of this behaviour in this theory. However, it is known from the analysis of Velizhanin:2022faj that the leading eigenvalue is not simply related to the one of $\mathcal{N}$$=$$4$ SYM. It would be interesting to study analytically the BFKL-type regime, where we take a double scaling limit of small $\omega\equiv S_{sl(2)}+1$ and small $h$ with $h^{2}/\omega$ finite. The shape of the curve in this limit should be describable analytically with the QSC, in particular it would be interesting to generalize to ABJM the recent method of Ekhammar:2024neh to obtain several terms in the expansion. We leave this for future work.
@@
We also estimated some coefficients in the expansion of the Pomeron eigenvalue at weak coupling,
$\mathcal{S}_{0}=1+\sum_{n>0}\mathcal{S}_{0}^{(2n)}h^{2n},$ (188)
where fitting our numerical data we estimate $\mathcal{S}_{0}^{(2)}\simeq 26.6609(6)$ and $\mathcal{S}_{0}^{(4)}\simeq-2.11(5)\times 10^{3}$.
@@ 4.4 Trajectory interpolating odd spins
It is well-known that the analytic continuations interpolating even and odd values of the spin are distinct. We studied the trajectory interpolating the dimensions of the operators (186) but with $S_{sl(2)}\in 2\mathbb{N}+1$.
@@
Figure 9: Trajectory interpolating operators (186) with odd spins ($sl(2)$-like grading). The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black bullets, their shadow by circles. The weak coupling limit (indicated by the dotted red piecewise line) is in this case a sharp intersection of two linear trajectories.
@@
$\mathcal{S}_{0}^{\text{odd}}=\frac{1}{2}+\sum_{n>0}\mathcal{S}_{0}^{\text{odd}\,(2n)}h^{2n},$ (189)
we estimate $\mathcal{S}_{0}^{\text{odd}\,(2)}\simeq 21.09289(9)$ and $\mathcal{S}_{0}^{\text{odd}\,(4)}\simeq-0.1496(1)\times 10^{4}$.
@@
Let us observe that also on the odd-spin trajectory there is a second BPS protected point, where the anomalous dimension vanishes, namely the point $\Delta_{sl(2)}=0$, $S_{sl(2)}=-1$. From the point of view of the QSC, the special nature of this point is signalled by the vanishing of some of the coefficients in front of $\mathbf{P}$ functions, indeed, due to (200), we have
$\mathcal{A}_{12}\mathcal{A}_{34}(\Delta_{sl(2)}=0,S_{sl(2)}=-1)&=&0,\\ \mathcal{A}_{13}\mathcal{A}_{24}(\Delta_{sl(2)}=0,S_{sl(2)}=-1)&=&0,\\ \mathcal{A}_{14}\mathcal{A}_{23}(\Delta_{sl(2)}=0,S_{sl(2)}=-1)&=&1,\end{aligned}\quad\leftrightarrow\quad\begin{aligned} \mathcal{A}_{12}\mathcal{A}_{34}\Big{|}_{\text{BPS point}}&=&0,\\ \mathcal{A}_{13}\mathcal{A}_{24}\Big{|}_{\text{BPS point}}&=&0,\\ \mathcal{A}_{14}\mathcal{A}_{23}\Big{|}_{\text{BPS point}}&=&1.\end{aligned}$ (190)
This is the same behaviour occurring for the BPS point living on the even-spin trajectories, i.e. the one corresponding to the stress-energy tensor, and it is the way the QSC signals a shortening condition. The vanishing of the anomalous dimension at this point can be verified at 6 loops using the analytic continuation of the perturbative results in Velizhanin:2022faj , and is confirmed by our numerics.
@@ 4.5 Trajectories and Bridges connecting them: discovering the symmetry
$\mathcal{G}_{\text{Bridge}}\cong(\mathcal{G}_{\text{Regge}})^{T},$ (191)
@@
$\texttt{W}_{S}:\quad S_{\text{top}}\rightarrow-1-S_{\text{top}},\quad\text{ i.e. }\quad S_{\text{bottom}}\rightarrow 5-S_{\text{bottom}},\quad\text{ i.e. }\quad S_{sl(2)}\rightarrow 1-S_{sl(2)},$ (192)
@@
$\tau_{\text{top}}\rightarrow\tau_{\text{top}}+1+2S_{\text{top}},\quad\tau_{sl(2)}\rightarrow\tau_{sl(2)}-1+2S_{sl(2)},\quad\tau_{\text{bottom}}\rightarrow\tau_{\text{bottom}}-5+2S_{\text{bottom}}.$ (193)
Figure 11: This figure reproduces figure 1, but giving the $sl(2)$-like grading representatives of the trajectories in accordance with the present discussion. The coupling is $h=0.1$. The ($sl(2)$-grading copy of) the leading twist-$1$ trajectory is shown, together with two Bridging trajectories connecting it to a subleading sheet with twist $5$. The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black bullets, shadows of local operators below unitarity by circles.
Figure 12: The panel shows, at coupling $h=0.1$ ($sl(2)$-like grading), the twist-$1$ trajectory interpolating between odd $S_{sl(2)}$ states, together with two Bridging trajectories connecting it to a subleading sheet with twist $5$. The unitarity-allowed region is depicted in gray. Physical states belonging to the spectrum are identified by black bullets, their shadow by circles.
@@
The name is explained if we follow the Bridging trajectories in the $(S,\Delta)$ plane. As we move $\Delta$, they will cross various integer values of the spin, and for each of these points we can determine whether they correspond to local operators by monitoring the gluing matrix, according to the locality criterion of Section 2.3. We find that, as the spin decreases in steps of two, they invariably reach local operators (or their spin-shadows) on subleading Regge trajectories, with twist increased in steps of four. At the points of intersection, we can reconstruct the full subleading trajectory by switching to the standard form of gluing matrix, $\mathcal{G}_{\text{Regge}}$. In this way, we can reach subleading sheets of the spectral surface while keeping our equations totally real.
For example, in Figure 11 is shown the Bridge shooting off from the $S_{sl(2)}=2$ operator on the leading Regge trajectory. We find that at $S_{sl(2)}=0$ (which is below the unitarity bound), it intersects a subleading Regge trajectory with twist $5$. The Figure also shows a second Bridge between the same two trajectories, the one which intersects the local operator with $S_{sl(2)}=4$ on the leading trajectory and which reaches the same subleading trajectory at a physical local operator with $S_{sl(2)}=2$. It is obvious to guess that there are in fact infinitely many Bridges between these two Trajectories, one for each non-protected local operator on the first trajectory. We have verified this conjecture in several experiments, finding Bridges passing through every non-BPS local operator we have inspected. [^33] Notice that there is instead no Bridge passing through the BPS point. If we continue the Bridges further increasing $\Delta$ past the twist-$5$ subleading trajectory, we find that they allow us to reach further subleading sheets, with twist always increasing in steps of four.
We tested these numerical findings for various values of the coupling, in particular $h=0.1$ and $h=0.27$. We believe that the structure is general, and that in this way it is possible to reach infinitely many subleading Regge trajectories by using alternatively the two forms of the gluing matrix.
@@
For instance, in Figure 11 we see a Bridge passing through the $S_{sl(2)}=2$ operator on the leading twist-$1$ Regge trajectory. This can also be seen as the spin-flip of a trajectory with twist 4. The Bridge intersecting the next operator with $S_{sl(2)}=4$ is a spin-flip of a trajectory with twist $8$, and so on: subsequent Bridges, once flipped, will be seen as trajectories of increasing twists in steps of four, as per (193).
Twist $\leftrightarrow$ co-twist symmetry.
The consequence of this observation is that in the Chew-Frautschi plot (which contains only Regge trajectories, not their spin-flips), we will have an unexpected symmetry relating points on different Regge trajectories. For each physical local operator sitting on a Regge trajectory of twist $\tau$, we will have an exact image, with the same anomalous dimension, which will sit below the unitarity bound on a subleading trajectory. The image is obtained by applying the Weyl reflection $\texttt{W}_{S}$, and thus where the original operator sits on a trajectory of twist $\tau$ and has co-twist $\tilde{\tau}$, its image will live on a subleading trajectory of twist $\tilde{\tau}$ and have co-twist $\tau$. On this second trajectory, the original operator (or rather its image) will appear below the unitarity bound. Thus we are learning something unexpected on the behaviour of trajectories below the unitarity bound, which is a less accessible region where the Lorentzian inversion formula does not apply. To make a very concrete example, the operator $S_{sl(2)}=2$ on the leading twist-$1$ trajectory will have an image below unitarity on a subleading trajectory with twist $4$.